A wheel is rotating at 30 revolutions per minute How many degrees has it rotated in 0.5 seconds?
1
Expert's answer
2019-05-08T09:42:38-0400
"1 \ rev=360 \degree"

"\theta =\omega t =\frac{30}{60} 0.5=0.25\ rev"

"\theta =0.25(360)=90\degree"